ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 167, 2008

Athanasiadis, C. A. and E. Tzanaki Shellability and higher Cohen-Macaulay connectivity of generalized cluster complexes

177

Bachurin, P., P. Bálint and I. P. Tóth Local ergodicity for systems with growth properties including multi-dimensional dispersing billiards

155

Bálint, P. See P. Bachurin

Berend, D. and M. D. Boshernitzan Nilpotent groups are round  49

Boshernitzan, M. D. See D. Berend

Button, J. O. Large groups of deficiency 1  111

Dutta, S. and V. P. Fonf On tree characterizations of G_{δ}- embeddings and some Banach spaces

27

Fonf, V. P. See S. Dutta

Fuchs, L. and R. Göbel Modules with absolute endomorphism rings  91
Gelander, T. On deformations of F_{n} in compact Lie groups  15

Göbel R. See L. Fuchs

Hai, Phùng Hô Tannaka-Krein duality for Hopf algebroids  193

Herzog, I. Contravariant functors on the category of finitely presented modules

347

Junge, M. and Quanhua Xu Noncommutative Burkholder/ Rosenthal inequalities II: applications

227

Kim, Joonil Maximal average along variable lines  1

León-Saavedra, F. and A. Piqueras-Lerena On weak positive supercyclicity

303

Libman, A. Webb’s conjecture for fusion systems  141
López, R. Rotational linear Weingarten surfaces of hyperbolic type  283

Muhly, P. S., D. Pask and M. Tomforde Strong shift equivalence of C^{∗}-correspondences

315

Pask, D. See P. S. Muhly

Pichot, M. Harmonic analysis from quasi-periodic domains  63

Piqueras-Lerena, A. See F. León-Saavedra Tomforde, M. See P. S. Muhly Tóth, I. P. See P. Bachurin Tzanaki, E. See C. A. Athanasiadis Xu, Quanhua See M. Junge